a father is now 24 years older than his son. In 8 years the father will be twice as old as his son will be then. Find there present ages
Answer by venugopalramana(3286)

a father is now 24 years older than his son. In 8 years the father will be twice as old as his son will be then. Find there present ages
let the son's age now be s
24 years older than this =24+s=age of father now
in 8 years son will be 8+s...twice as old as this =2*(8+s)=16+2s....(1)
in 8 years father will be 24+s+8=32+s.......(2)
it is given that (1)=(2).hence
16+2s=32+s
2s-s=32-16=16
s=16...hence son's present age is 16 years and father's present age is 24+16=40 years.